OpenStudy (anonymous):

Solve 3(2x – 4) + 2 = 14 and justify each step with a property. Be sure to show each step you do to solve it. You should use a total of 3 properties.

jigglypuff314 (jigglypuff314):

3(2x – 4) + 2 = 14 distributive property addition property of equality division property of equality x =

jigglypuff314 (jigglypuff314):

3(2x – 4) + 2 = 14 6x -12 + 2 = 14 6x - 10 = 14 +10 +10 6x = 24 /6 /6 x = ...

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

24 @jigglypuff314

OpenStudy (anonymous):

x = 4